# Continuous Ensemble Weather Forecasting with Diffusion models

### v1

@@ Appendix C Additional Results
An alternative to directly producing forecasts at a fine temporal resolution would be to simply interpolate the forecasts sampled using an autoregressive model. We argue that our ARCI method produces more accurate and realistic predictions than simple linear interpolation. Figure 9 shows the RMSE and Spread/Skill for a linear interpolation of AR-24h, compared to ARCI-24/6h. The interpolated forecasts behave much worse than the ARCI model on the same resolution, which strengthens the role of our model as an advanced interpolator.
@@
Table 6 and figure 10 show the results on 10 day forecasting for all variables. Figure 11 shows the results on forecasting with 1h resolution. Figure 12 shows the temporal difference for different values of $\rho$.
Table 6: Results on 5 and 10 day forecasting for all variables. For RMSE and CRPS, lower values are better, and SSR should be close to 1.
Lead time 5 days Lead time 10 days
Variable Model RMSE CRPS SSR RMSE CRPS SSR
z500 Deterministic 766.7 483.9 - 1042 661.5 -
Graph-EFM 699.1 317.5 1.13 817.1 373.6 1.1
AR-6h 602.3 287.8 0.75 811.8 391.9 0.88
AR-24h 544.2 242.7 0.84 750.6 335.2 0.94
CI-6h 707.8 321.2 0.59 885.7 406.6 0.6
ARCI-24/6h 560.9 256.7 0.86 765.6 355.2 0.93
t850 Deterministic 3.48 2.36 - 4.54 3.17 -
Graph-EFM 3.12 1.56 1.11 3.51 1.77 1.12
AR-6h 2.72 1.34 0.82 3.39 1.69 0.92
AR-24h 2.55 1.24 0.89 3.25 1.6 0.96
CI-6h 3.06 1.5 0.74 3.68 1.85 0.71
ARCI-24/6h 2.6 1.27 0.9 3.29 1.63 0.95
t2m Deterministic 2.71 1.72 - 3.56 2.28 -
Graph-EFM 2.51 1.1 1.09 2.88 1.32 1.14
AR-6h 2.13 0.96 0.83 2.62 1.18 0.89
AR-24h 1.98 0.87 0.9 2.48 1.09 0.95
CI-6h 2.29 1.0 0.79 2.75 1.21 0.75
ARCI-24/6h 2.02 0.9 0.9 2.51 1.11 0.94
u10 Deterministic 4.37 2.93 - 5.14 3.53 -
Graph-EFM 3.81 1.93 0.97 4.08 2.07 0.97
AR-6h 3.47 1.71 0.86 3.95 1.97 0.94
AR-24h 3.32 1.62 0.92 3.85 1.9 0.97
CI-6h 3.87 1.91 0.77 4.27 2.14 0.78
ARCI-24/6h 3.35 1.64 0.93 3.87 1.92 0.97
v10 Deterministic 4.48 3.0 - 5.23 3.58 -
Graph-EFM 3.88 1.96 0.94 4.11 2.08 0.93
AR-6h 3.55 1.76 0.86 4.02 2.0 0.96
AR-24h 3.42 1.68 0.93 3.96 1.96 0.99
CI-6h 4.0 1.99 0.77 4.39 2.2 0.8
ARCI-24/6h 3.45 1.69 0.94 3.97 1.97 0.99
ws10 Deterministic 3.09 2.2 - 3.44 2.49 -
Graph-EFM 2.56 1.35 1.01 2.65 1.4 1.01
AR-6h 2.38 1.23 0.9 2.57 1.34 0.96
AR-24h 2.3 1.18 0.96 2.53 1.31 0.99
CI-6h 2.54 1.32 0.88 2.68 1.4 0.89
ARCI-24/6h 2.31 1.19 0.96 2.54 1.32 0.99

### v2

@@ Appendix C Additional Results
An alternative to directly producing forecasts at a fine temporal resolution would be to simply interpolate the forecasts sampled using an autoregressive model. We argue that our ARCI method produces more accurate and realistic predictions than simple linear interpolation. Figure 8 shows the RMSE and Spread/Skill for a linear interpolation of AR-24h, compared to ARCI-24/6h. The interpolated forecasts behave much worse than the ARCI model on the same resolution, dismissing this argument.
@@
Table 6, 7 and show the results on 5 and 10 day forecasting for all variables. The error bars given for ARCI-24/6h is a standard deviation calculated by retraining the model five times and evaluating each model on 10 ensemble members. This shows that the model is robust to network initialization. Figure 9 show the results on forecasting at 6h resolution for all variables. Figure 10 shows the same results on forecasting with 1h resolution. Figure 11 shows the temporal difference for different values of $\rho$.
Table 6: Results on 5 day forecasting for all variables. For RMSE and CRPS, lower values are better, and SSR should be close to 1.
Lead time 5 days
Variable Model RMSE CRPS SSR
z500 Deterministic 766.7 483.9 -
Graph-EFM 699.1 317.5 1.13
AR-6h 602.3 287.8 0.75
AR-24h 544.2 242.7 0.84
CI-6h 707.8 321.2 0.59
ARCI-24/6h 560.9 $\pm$ 15.6 256.7 $\pm$ 20.2 0.86 $\pm$ 0.015
t850 Deterministic 3.48 2.36 -
Graph-EFM 3.12 1.56 1.11
AR-6h 2.72 1.34 0.82
AR-24h 2.55 1.24 0.89
CI-6h 3.06 1.5 0.74
ARCI-24/6h 2.6 $\pm$ 0.031 1.27 $\pm$ 0.021 0.9 $\pm$ 0.012
t2m Deterministic 2.71 1.72 -
Graph-EFM 2.51 1.1 1.09
AR-6h 2.13 0.96 0.83
AR-24h 1.98 0.87 0.9
CI-6h 2.29 1.0 0.79
ARCI-24/6h 2.02 $\pm$ 0.036 0.9 $\pm$ 0.035 0.9 $\pm$ 0.012
u10 Deterministic 4.37 2.93 -
Graph-EFM 3.81 1.93 0.97
AR-6h 3.47 1.71 0.86
AR-24h 3.32 1.62 0.92
CI-6h 3.87 1.91 0.77
ARCI-24/6h 3.35 $\pm$ 0.032 1.64 $\pm$ 0.019 0.93 $\pm$ 0.007
v10 Deterministic 4.48 3.0 -
Graph-EFM 3.88 1.96 0.94
AR-6h 3.55 1.76 0.86
AR-24h 3.42 1.68 0.93
CI-6h 4.0 1.99 0.77
ARCI-24/6h 3.45 $\pm$ 0.026 1.69 $\pm$ 0.014 0.94 $\pm$ 0.006
ws10 Deterministic 3.09 2.2 -
Graph-EFM 2.56 1.35 1.01
AR-6h 2.38 1.23 0.9
AR-24h 2.3 1.18 0.96
CI-6h 2.54 1.32 0.88
ARCI-24/6h 2.31 $\pm$ 0.016 1.19 $\pm$ 0.01 0.96 $\pm$ 0.006
Table 7: Results on 10 day forecasting for all variables. For RMSE and CRPS, lower values are better, and SSR should be close to 1.
Lead time 10 days
Variable Model RMSE CRPS SSR
z500 Deterministic 1042 661.5 -
Graph-EFM 817.1 373.6 1.1
AR-6h 811.8 391.9 0.88
AR-24h 750.6 335.2 0.94
CI-6h 885.7 406.6 0.6
ARCI-24/6h 765.6 $\pm$ 21.4 355.2 $\pm$ 33.0 0.93 $\pm$ 0.018
t850 Deterministic 4.54 3.17 -
Graph-EFM 3.51 1.77 1.12
AR-6h 3.39 1.69 0.92
AR-24h 3.25 1.6 0.96
CI-6h 3.68 1.85 0.71
ARCI-24/6h 3.29 $\pm$ 0.05 1.63 $\pm$ 0.041 0.95 $\pm$ 0.007
t2m Deterministic 3.56 2.28 -
Graph-EFM 2.88 1.32 1.14
AR-6h 2.62 1.18 0.89
AR-24h 2.48 1.09 0.95
CI-6h 2.75 1.21 0.75
ARCI-24/6h 2.51 $\pm$ 0.068 1.11 $\pm$ 0.076 0.94 $\pm$ 0.017
u10 Deterministic 5.14 3.53 -
Graph-EFM 4.08 2.07 0.97
AR-6h 3.95 1.97 0.94
AR-24h 3.85 1.9 0.97
CI-6h 4.27 2.14 0.78
ARCI-24/6h 3.87 $\pm$ 0.03 1.92 $\pm$ 0.018 0.97 $\pm$ 0.01
v10 Deterministic 5.23 3.58 -
Graph-EFM 4.11 2.08 0.93
AR-6h 4.02 2.0 0.96
AR-24h 3.96 1.96 0.99
CI-6h 4.39 2.2 0.8
ARCI-24/6h 3.97 $\pm$ 0.018 1.97 $\pm$ 0.011 0.99 $\pm$ 0.01
ws10 Deterministic 3.44 2.49 -
Graph-EFM 2.65 1.4 1.01
AR-6h 2.57 1.34 0.96
AR-24h 2.53 1.31 0.99
CI-6h 2.68 1.4 0.89
ARCI-24/6h 2.54 $\pm$ 0.011 1.32 $\pm$ 0.008 0.99 $\pm$ 0.008
@@ Appendix D Example forecasts
Appendix E Continuity of solutions to Probability Flow ODE
The proposed method in section 4 generates continuous ensemble trajectories by introducing temporal correlations in the noise. Sampling is performed by solving the probability flow ODE:
$\mathrm{d}{z}(s)=-sS_{\theta}({z}(s),s;X[k{-}M{:}k],t)\mathrm{d}s,\quad{z}(1)={z}_{0},$ (8)
starting in pure noise $z_{0}\sim p_{\text{noise}}$ and ending in a forecast ${z}(0)\sim p(X[k{+}1]|X[k{-}M{:}k])$. In practice, the integration is only done up to some time $\epsilon\ll 1$ to overcome instability caused by the singularity of the score function, $\nabla\log p_{s}$, at $s=0$.
To ensure that the trajectories are continuous, we must show that solutions to eq. 8 are continuous functions of the lead time $t$. Here, we consider the case where the initial noise ${z}_{0}$ remains fixed across different $t$. The more general scenario, where ${z}_{0}$ follows a stochastic noise process, is left for future work.
Let $f(s,z,t):=-sS_{\theta}({z},s;X[k{-}M{:}k],t)$ defined for $(s,z)\in D:=[\epsilon,1]\times\mathbb{R}^{n}$ and $t\in[0,T]$. The dependence on the previous states $X[k{-}M{:}k]$ has been suppressed in the notation since all forecasts are conditioned on a fixed $X[k{-}M{:}k]$. To phrase the result, we first make an assumption on the regularity of $f$.
Assumption 1.
Let $f(s,z,t)$ be continuous in $(s,z,t)$ and locally Lipschitz in $z$, i.e. for any closed bounded set $U\subset D$ there is a $K_{U}$ such that
$|f(s,z_{1},t)-f(s,z_{2},t)|\leq K_{U}|z_{1}-z_{2}|$
for any $(s,z_{1}),(s,z_{2})\in U$ and $t\in[0,T]$.
If $f(s,z,t)$ is continuous for $(s,z)\in D$, then by the fundamental existence theorem for ODEs (Theorem 1.1 in Hale (2009)), there exists at least one solution of eq. 8 passing through any given point $(s_{0},z_{0})\in D$. Suppose further this solution is unique and denote it as $z(s,s_{0},z_{0},t)$. For any $(s_{0},z_{0})\in D$ and $t\in[0,T]$, let $(a(s_{0},z_{0},t),b(s_{0},z_{0},t))$ be the maximal interval of existence of $z(s,s_{0},z_{0},t)$. We define the set
$E:=\{(s,s_{0},z_{0},t):a(s_{0},z_{0},t)<s<b(s_{0},z_{0},t),(s_{0},z_{0})\in D,t\in[0,T]\}.$
as the domain of definition of $z(s,s_{0},z_{0},t)$. Since our primary interest lies in the continuity of solutions, we assume access to some domain of definition $E$ without explicitly constructing it. For a proof of global existence under stronger assumptions, see theorem B in Section 70, Chapter 13 of Simmons (2016). With this, we can state the main result.
Theorem 1 (Theorem 3.2 in Hale (2009)).
Assume that $f$ satisfies assumption 1 and consider the initial value problem
$\mathrm{d}{z}(s)=f(s,z,t)\mathrm{d}s,\quad{z}(s_{0})={z}_{0}.$ (9)
Then, for every $(s_{0},z_{0})\in D$ and $t\in[0,T]$, there is a unique solution $z(s,s_{0},z_{0},t)$ with $z(s_{0},s_{0},z_{0},t)=z_{0}$, and this solution is continuous in $(s,s_{0},z_{0},t)$ within its domain of definition.
Theorem 1 shows that the solutions to the probability flow ODE eq. 8 are continuous as functions of lead-time if $-sS_{\theta}({z},s;X[k{-}M{:}k],t)$ is continuous in $(z,s,t)$ and locally Lipschitz in ${z}$. Since the neural network $S_{\theta}$ is a composition of continuous functions, it is necessarily a continuous function. The remaining assumption on $S_{\theta}$ being locally Lipschitz in ${z}$ is not a particularly strong one, and commonly assumed when proving results about neural networks (Karras et al., 2022; Song et al., 2020; Albergo & Vanden-Eijnden, 2023). To calculate the Lipschitz constant one could for example use the method proposed in Virmaux & Scaman (2018).